Given the quadratic function f(x) = x2 – 12x + 36, find a value of x such that f(x) = 9.


Answer by edjones(8007)

x^2-12x+36=9
x^2-12x+27=0
(x-9)(x-3)=0
x=9, x=3
.
Ed